OpenStudy (anonymous):

how do I find the value of this expression? 12+{10/[11-3(2)]}

OpenStudy (anonymous):

first -3x2=-6 11-6=5 10/5=2 12+2=14

OpenStudy (beccaboo333):

Welcome to OpenStudy, next time try posting in the Mathematics section on here. Good work @FLVSenglishstudent